OpenStudy (anonymous):

Write the equation of the line with an undefined slope, passing through the point (2, 5)

ganeshie8 (ganeshie8):

hint : vertical line has undefined slope

OpenStudy (dan815):

in defined means its infinite slope or vertical

OpenStudy (dan815):

un*

OpenStudy (anonymous):

so m is 0?

ganeshie8 (ganeshie8):

undefined = 1/0

OpenStudy (anonymous):

y-y1 = m(x-x1)

OpenStudy (dan815):



OpenStudy (shamim):

slope\[m=\tan 90=\frac{ \sin 90 }{ \cos 90 }=\frac{ 1 }{ 0 }\]

OpenStudy (anonymous):

whats the equation? y - 5 = 0(x - 2) y = 5 is that right?

OpenStudy (shamim):

no

OpenStudy (shamim):

it will b\[y-5=\frac{ 1 }{ 0 }(x-2)\]

OpenStudy (anonymous):

no more simplification needed?

OpenStudy (shamim):

\[x-2=0\]

OpenStudy (anonymous):

but u deleted it woman. does not count.

OpenStudy (anonymous):

u should not have deleted it. i would of given u the medal if u had left it, your confidence must be down. believe in yourself a bit more.

OpenStudy (anonymous):

i wish it would let me give out more medals on here thnaks everyone.

OpenStudy (anonymous):

bit harsh, sorry i did not give u the medal. i will have some more questions later, maybe u can help me again.